# POGEMA: A Benchmark Platform for Cooperative Multi-Agent Navigation

### v1

@@ B.6 Pathfinding
Algorithm Makespan
DCC $189.56$ p m 28.28
IQL $1825.95$ p m 137.11
LaCAM $179.82$ p m 20.21
MAMBA $416.45$ p m 139.34
QMIX $955.54$ p m 203.76
QPLEX $933.74$ p m 204.21
SCRIMP $1460.04$ p m 176.27
VDN $1733.20$ p m 158.91
To compute Pathfidning metric we run the approaches on the instances with a single agent. For this purpose we utilized large MovingAI mapf with $256\times 256$ size, the results are presented in Table 6. While this task seems easy, most of the hybrid and MARL approaches are not able to effectively solve them. Only LaCAM is able to find optimal paths in all the cases, as it utilizes precomputed costs to the goal location as a heuristic. Most of the evaluated hybrid and MARL approaches are also contain a sort of global guidance in one the channels of their observations. However, large maps with out-of-distribution structure, the absence of communication and other agents in local observations are able to lead to inconsistent behavior of the models that are not able to effectively choose the actions that lead the agent to its goal. Please note, SoC and makespan metrics in this case are equal, as there is only one agent in every instance.
@@ Appendix C Results for LifeLong MAPF Benchmark
In this section, we present the extended results of the LifeLong MAPF benchmark analysis, highlighting performance, out-of-distribution handling, scalability, cooperation, congestion, and pathfinding.
@@ C.1 Performance
Performance metric in LMAPF case is based on the ratio of throughput compared to the best obtained one. In contrast to SoC, throughput should be as high as possible. There is also no CSR metric, as there is no need for agents to be at their goal locations simultaneously. As well as in MAPF case, the best results are obtained by centralized search-based approach – RHCR. The best results among decentralized methods demonstrate Follower and MATS-LP, Following them, comparable results are shown by ASwitcher and LSwitcher. Between MARL methods the highest throughput on both Random and Mazes maps is obtained by MAMBA.
@@ C.2 Out-of-Distribution
The evaluation on out-of-distribution set of maps confirms the results obtained on Random and Mazes maps. The best results demonstrates RHCR. Next best results are obtained by Follower and MATS-LP, which are much closer to RHCR in this experiment. While MATS-LP outperforms Follower on the instances with 64, 128 and 192 agents, Follower is still better on the instances with 256 agents. Such relation is probably explained by the presence of dynamic edge-costs in Follower that allows to better distribute agents along the map and reduce congestion between them.
@@ C.4 Cooperation
VDN $0.030$ p m 0.004
As well as for MAPF setting, cooperation metric is computed based on the results obtained on Puzzles dataset. Table 8 contains average throughput obtained by each of the evaluated approaches. Here again the best results are obtained by RHCR algorithm. In contrast to Random, Mazes and Warehouse sets of maps, where MATS-LP and Follower demonstrate close results, the ability to simulate the behavior of other agents, provided by MCTS in MATS-LP, allows to significantly outperform Follower on small Puzzles maps. The rest approaches demonstrate much worse results, especially IQL, QPLEX and VDN that have 10 times worse average throughput than RHCR.
C.5 Congestion
Table 9 contains average agents density presented in observations. As it was already mentioned in LMAPF:Out-of-Distribution section, Follower contains a mechanism that allows to effectively distribute agents along the map. As a result, the lowest density is demonstrated exactly by this approach. MARL methods, such as MAMBA and QMIX are also demonstrate low average agents density, as they actually utilize the same observation as Follower does.
Table 9: Average Agents Density by Number of Agents
Algorithm 64 Agents 96 Agents 128 Agents 160 Agents 192 Agents
ASwitcher $0.074$ p m 0.001 $0.111$ p m 0.001 $0.146$ p m 0.001 $0.176$ p m 0.001 $0.203$ p m 0.001
Follower $0.073$ p m 0.000 $0.101$ p m 0.000 $0.126$ p m 0.000 $0.150$ p m 0.000 $0.173$ p m 0.000
IQL $0.080$ p m 0.001 $0.127$ p m 0.002 $0.188$ p m 0.003 $0.257$ p m 0.003 $0.319$ p m 0.002
LSwitcher $0.075$ p m 0.001 $0.114$ p m 0.001 $0.149$ p m 0.001 $0.180$ p m 0.001 $0.208$ p m 0.001
MAMBA $0.073$ p m 0.000 $0.095$ p m 0.000 $0.119$ p m 0.001 $0.146$ p m 0.001 $0.176$ p m 0.001
MATS-LP $0.110$ p m 0.002 $0.176$ p m 0.002 $0.231$ p m 0.003 $0.274$ p m 0.003 $0.306$ p m 0.003
QMIX $0.074$ p m 0.000 $0.100$ p m 0.000 $0.126$ p m 0.000 $0.150$ p m 0.001 $0.175$ p m 0.001
QPLEX $0.078$ p m 0.000 $0.114$ p m 0.001 $0.147$ p m 0.001 $0.176$ p m 0.001 $0.216$ p m 0.002
RHCR $0.088$ p m 0.000 $0.125$ p m 0.001 $0.170$ p m 0.002 $0.242$ p m 0.004 $0.314$ p m 0.004
VDN $0.077$ p m 0.001 $0.109$ p m 0.001 $0.141$ p m 0.001 $0.173$ p m 0.001 $0.204$ p m 0.002
C.6 Pathfinding
Table 10: Pathfinding results.
@@ C.6 Pathfinding
VDN $1733.20$ p m 157.96
Pathfinding metric is tailored to indicate how well the algorithm is able to guide am agent to its goal location. As a result, there is actually no need to run the algorithms on LifeLong instances. Instead, they were run on the same set of instances that were utilized for MAPF approaches. The results of this evaluation are presented in Table 10. Again, the best results were obtained by search-based approach – RHCR. Its implementation was slightly modified to work on MAPF instances, when there is no new goal after reaching the current one. Either optimal or close to optimal paths are able to find Follower and MATS-LP. Followers misses optimal paths due to the integrated technique that changes the edge-costs. MATS-LP adds noise to the root of the search tree that might result in choosing of wrong actions. For the approaches from Switcher family it’s actually almost impossible to find optimal paths as they have no information about global map and operate only based on the local observations.
@@ Appendix D Code examples for POGEMA
{listing}
[…]
POGEMA is an environment that provides a simple scheme for creating MAPF scenarios, specifying the parameters of GridConfig. The main parameters are: on_target (the behavior of an agent on the target, e.g., restart for LifeLong MAPF and nothing for classical MAPF), seed – to preserve the same generation of the map; agent; and their targets for scenario, size – used for cases without custom maps to specify the size of the map, density – the density of obstacles, num_agents – the number of agents, obs_radius – observation radius, collision_system – controls how conflicts are handled in the environment (we used a soft collision system for all of our experiments). The example of creation such instance is presenten in Listing D.
Visualization of the agents is a crucial tool for debugging algorithms, visually comparing them, and presenting the results. Many existing MARL environments lack such tools, or have limited visualization functionality, e.g., requiring running the simulator to provide replays, or offering visualizations only in one format (such as videos). In the POGEMA environment, there are three types of visualization formats. The first one is console rendering, which can be used with the default render methods of the environment; this approach is useful for local or server-side debugging. The preferred second option is SVG animations. An example of generating such a visualization is presented in the listing above. This approach allows displaying the results using any browser. It provides the ability to highlight high-quality static images (e.g., as the images provided in the paper) or to display results on a website (e. g., animations of the POGEMA repository on GitHub). This format ensures high-quality vector graphics. The third option is to render the results to video format, which is useful for presentations and videos.
@@ Appendix E POGEMA Toolbox
The first one is registries to handle custom maps and algorithms. To create a custom map, the user first needs to define it using ASCII symbols or by uploading it from a file, and then register it using the toolbox (see Listing E). The same approach is used to register and create algorithms (see Listing E). In that listing, the registration of a simple algorithm is presented, which must includ two methods: act and reset_states. This approach can also accommodate a set of hyperparameters which the Toolbox handles.
{listing}
[…]
{listing}
[…]
{listing}
[…]
Second, it provides a unified way of conducting distributed testing using Dask  [^11] https://github.com/dask/dask and defined configurations. An example of such a configuration is provided in Listing E. The configuration is split into three main sections; the first one details the parameters of the POGEMA environment used for testing. It also includes iteration over the number of agents, seeds, and names of the map (which were registered beforehand). The unified grid_search tag allows for the examination of any existing parameter of the environment. The second part of the configurations is a list of algorithms to be tested. Each algorithm has its alias (which will be shown in the results) and name, which specifies the family of methods. It also includes a list of hyperparameters common to different approaches, e.g., number of processes, parallel backend, etc., and the specific parameters of the algorithm.
@@ Appendix F Extended Related Work
Nevertheless, some of these drawbacks have already been addressed. SMAX [39] provides a hardware-accelerated JAX version of the environment, but it cannot guarantee full compatibility since the StarCraft II engine is proprietary software. SMAC v2 [13] solves the problem of determinism, highlighting this issue in the original SMAC environments. Moreover, an evaluation protocol for the SMAC environment is proposed in [15]. Despite these efforts, it’s hard to say that these tasks require the generalization ability of the agent, since the training and evaluation are conducted on the same scenario.
@@
Multi-agent MuJoCo — In MAMuJoCo, the standard tasks involve agents controlling different sets of joints (or a single joint) within a simulated robot. This set of environments is a natural adaptation of the environment presented in the well-known MuJoCo physics engine [55]. These tasks don’t require high generalization abilities or navigation. In the newer version, MuJoCo provides a hardware-accelerated version, forming the basis for Multi-agent BRAX [39], which enhances performance and efficiency.
@@
Google Research Football — Google Research Football [21] is a multi-agent football simulator that provides a framework for cooperative or competitive multi-agent tasks. Despite the large number of possible scenarios in the football academy and the requirement for simple navigation, the tasks are highly specific to the studied domain. Additionally, the number of possible agents is limited. Moreover, the framework offers low scalability, requiring a heavy engine.
@@
Multi-robot warehouse — The multi-robot warehouse environment RWARE [35] simulates a warehouse with robots delivering requested goods. The environment is highly specific to delivery tasks; however, it doesn’t support procedurally generated scenarios, thus not requiring generalization abilities or an evaluation protocol. The best-performing solution [8] in this environment is trained on only 4 agents. The benchmark is highly related to multi-agent pathfinding tasks; however, it doesn’t provide centralized solution integration, which could serve both as an upper bound for learnable decentralized methods and as a source of expert demonstrations.
@@
Level-Based Foraging — Multi-agent environment LBF [35] simulates food collection by several autonomously navigating agants in a grid world. Each agent is assigned a level. Food is also randomly scattered, each having a level on its own. The collection of food is successful only if the sum of the levels of the agents involved in loading is equal to or higher than the level of the food. The agents are getting rewarded by level of food they collected normalized by their level and overall food level of the episode. The game requires cooperation but also the agents can emerge competitve behavior. The environment is very efficiently designed and very simple to set up; however, it doesn’t support procedurally generated scenarios, thus not requiring generalization abilities or an evaluation protocol.
@@
Flatland — The Flatland environment [30] is designed to address the specific problem of fast, conflict-free train scheduling on a fixed railway map. This environment was created for the Flatland Competition [22]. The overall task is centralized with full observation; however, there is an adaptation to partial observability for RL agents. Unfortunately, during several competitions, despite the presence of stochastic events, centralized solutions [24] from operations research field have outperformed RL solutions by a large margin in both quality and speed. The environment is procedurally generated, which requires high generalization abilities, and the benchmark provides an evaluation protocol. A significant drawback is the extremely slow speed of the environment, which highly restricts large-scale learning.
@@
Overcooked — The Overcooked is a benchmark environment for fully cooperative human-AI task performance, based on the widely popular video game [7]. In the game, agents control chefs tasked to cook some dishes. Due to possible complexity of the cooking process, involving multistep decision-making, it requires emergence of cooperative behaviour between the agents.
@@
Griddly — This is a grid-based game engine [4], allowing to make various and diverse grid-world scenarios. The environment is very performance efficient, being able to make thousands step per second. Moreover, there is test coverage and continuous integration support, allowing open-source development. The engine provides support for different observation setups and maintains state history, making it useful for search based methods.
@@
Multi-player game simulators — Despite the popularity of multi-player games, it’s a challenging problem to develop simulators of the games that could be used for research purposes. One of the most popular adaptations are MineCraft MALMO [19] that allows to utilise MineCraft as a configurable research platform for multi-agent research and model various agents’ interactions. In spite of the game’s flexible functionality, it depends on external runtime, so might be very hard to set up or extremely slow to iterate with. That’s why there are several alternatives that prioritise fast iteration over the environment complexity, like Neural MMO [51] that models a simple MMO RPG with agents with a shared resource pool. On top of that, there are even faster implementation, targeting coordination or cooperation, like Hide-and-Seek [3], modelling competition, or GoBigger [69], focusing on competition between cooperating populations.
@@
Multi-agent Driving Simulators — Autonomous driving is one of the important practical applications of MARL, and Nocturne  [57] is a 2D simulator, written in C++, that focuses on different scenarios of interactions — e.g. intersections, roundabouts etc. The simulator is based on trajectories collected in real life, so allows modelling practical scenarios. This environment has evaluation protocols and supports open-source development with continuous integration and covered by tests. There are also environments, focusing on particular details of driving, for example, Multi-car Racing  [44] that represents racing from bird’s eye view.
@@
Suits of multi-agent environments — These multi-agent environments are designed to be very simple benchmarks for specific tasks. Jumanji  [5] is a set of environments for different multi-agent scenarios connected to combinatorial optimization and control, for example, routing or packing problems. With the purpose for each environment to be focused on the particular task, the overall suit doesn’t test generalization or enable procedural generation. Multi Particle Environments (MPE)  [28] is a communication oriented set of partially observable environments where particle agents are able to interact with fixed landmarks and each other, communicating with each other. SISL  [16] is a set of three dense reward environments was developed to have a simple benchmark for various cooperative scenarios. For environment suits, testing generalization, MeltingPot  [1] comes into place. This set of the environments contains a diverse set of cooperative and general-sum partially observable games and maintains two populations of agents: focal (learning) and visiting (unknown to the environment) to benchmark generalization abilities of MARL algorithms. The set in based on the own game engine and might be extended quite easily.
@@
Real-world Engineering in Practice — Real-world engineering tasks can often be addressed by sophisticated MARL solutions. IMP-MARL [23] provides a platform for evaluating the scalability of cooperative MARL methods responsible for planning inspections and repairs for specific system components, with the goal of minimizing maintenance costs. At the same time, agents must cooperate to minimize the overall risk of system failure. MATE [34] addresses target coverage control challenges in real-world scenarios. It presents an asymmetric cooperative-competitive game featuring two groups of learning agents, cameras and targets, each with opposing goals.

### v2

@@ B.5 Pathfinding
To compute Pathfinding metric we run the approaches on the instances with a single agent. For this purpose we utilized large Cities maps with $256\times 256$ size, the results are presented in Table 5. While this task seems easy, most of the hybrid and MARL approaches are not able to effectively solve them. Only LaCAM is able to find optimal paths in all the cases, as it utilizes precomputed costs to the goal location as a heuristic. Most of the evaluated hybrid and MARL approaches are also contain a sort of global guidance in one the channels of their observations. However, large maps with out-of-distribution structure, the absence of communication and other agents in local observations are able to lead to inconsistent behavior of the models that are not able to effectively choose the actions that lead the agent to its goal. Please note, SoC and makespan metrics in this case are equal, as there is only one agent in every instance.
B.6 Coordination
The coordination metric is based on the number of collisions that occur during an episode. These collisions can occur either between agents, when two or more agents try to occupy the same cell or traverse the same edge simultaneously, or with static obstacles, when an agent tries to occupy a blocked cell. All such collisions are prevented by POGEMA during the action execution process, as all colliding actions are replaced with waiting actions instead. Figure 7 shows the average total number of collisions that occurred while solving instances with the corresponding number of agents on the Mazes and Random map sets.
Figure 7: Total number of collisions occurred during solving MAPF instances with corresponding number of agents of Mazes and Random sets of maps.
The highest number of collisions occurred with the MAMBA and VDN approaches, while the fewest were with DCC. The low number of collisions made by DCC compared to other approaches can be attributed to the presence of a communication mechanism that helps avoid collisions. The results of two evaluated approaches, LaCAM and SCRIMP, were omitted and are not presented in Figure 7. LaCAM is a centralized planner, so its solutions are collision-free by design. SCRIMP has an integrated environment with communication and tie-breaking mechanisms that resolve all collisions.
@@ Appendix C Results for LifeLong MAPF Benchmark
In this section, we present the extended results of the LifeLong MAPF benchmark analysis, highlighting performance, out-of-distribution handling, scalability, cooperation, coordination, and pathfinding.
@@ C.1 Performance
Performance metric in LMAPF case is based on the ratio of throughput compared to the best obtained one. In contrast to SoC, throughput should be as high as possible. There is also no CSR metric, as there is no need for agents to be at their goal locations simultaneously. As well as in MAPF case, the best results are obtained by centralized search-based approach – RHCR. The best results among decentralized methods demonstrate Follower and MATS-LP, following them, comparable results are shown by QPLEX, QMIX, ASwitcher which significantly outperforms MAMBA on both Mazes and Random maps. The results are presented in Figure 8.
@@ C.2 Out-of-Distribution
The evaluation on out-of-distribution set of maps confirms the results obtained on Random and Mazes maps. The best results demonstrates RHCR. Next best results are obtained by Follower and MATS-LP, which are much closer to RHCR in this experiment. While MATS-LP outperforms Follower on the instances with 64, 128 and 192 agents, Follower is still better on the instances with 256 agents. Such relation is probably explained by the presence of dynamic edge-costs in Follower that allows to better distribute agents along the map and reduce coordination between them.
@@ C.4 Cooperation
VDN $0.144$ p m 0.015
As well as for MAPF setting, cooperation metric is computed based on the results obtained on Puzzles dataset. Table 7 contains average throughput obtained by each of the evaluated approaches. Here again the best results are obtained by RHCR algorithm. In contrast to Random, Mazes and Warehouse sets of maps, where MATS-LP and Follower demonstrate close results, the ability to simulate the behavior of other agents, provided by MCTS in MATS-LP, allows to significantly outperform Follower on small Puzzles maps. The rest approaches demonstrate much worse results, especially IQL, MAMBA, VDN that have almost 5 times worse average throughput than RHCR.
C.5 Pathfinding
@@
VDN $1609.50$ p m 172.46
Table 8: Pathfinding results.
Pathfinding metric is tailored to indicate how well the algorithm is able to guide an agent to its goal location. As a result, there is actually no need to run the algorithms on LifeLong instances. Instead, they were run on the same set of instances that were utilized for MAPF approaches.
The results of this evaluation are presented in Table 8. Again, the best results were obtained by search-based approach – RHCR. Its implementation was slightly modified to work on MAPF instances, when there is no new goal after reaching the current one. Either optimal or close to optimal paths are able to find MATS-LP, Follower and QPLEX. Followers misses optimal paths due to the integrated technique that changes the edge-costs. MATS-LP adds noise to the root of the search tree that might result in choosing of wrong actions. For approaches in the Switcher family, it is nearly impossible to find optimal paths, as they lack information about the global map and rely solely on local observations. Surprisingly, ASwitcher outperforms MAMBA, QMIX, IQL, and VDN, which are provided with a global map.
C.6 Coordination
Figure 10 illustrates the average total number of collisions that occurred during the solving of LMAPF instances. The absolute values are higher than those obtained during the solving of MAPF instances. This behavior is explained by the extended episode length in LMAPF instances, which is twice as long. Moreover, in MAPF scenarios, the episode can end when all agents reach their goal locations, whereas in LMAPF scenarios, all agents continue to act until the episode length limit is reached.
All MARL approaches show poor results, with MAMBA being the worst among them. The fewest collisions are made by the Switcher approaches, i.e., ASwitcher and LSwitcher. A comparable number of collisions is demonstrated by MATS-LP and Follower on the Random set of maps. The difference in the behavior of these two approaches on the Random and Mazes map sets is likely due to the more complex structure of obstacles on the Mazes maps, where their heuristic guidance more often leads to collisions. It’s also worth noting that MATS-LP has no collisions with static obstacles, as it employs a masking mechanism that prevents selecting an action that leads an agent into a blocked cell. Such a mechanism could be implemented in other approaches to prevent this type of collision and potentially improve their performance. The results of the RHCR approach are omitted, as it is a centralized planner and its solutions are guaranteed to be collision-free.
Figure 10: Total number of collisions occurred during solving LMAPF instances with corresponding number of agents of Mazes and Random sets of maps.
@@ Appendix D Code Examples for POGEMA
POGEMA is an environment that provides a simple scheme for creating MAPF scenarios, specifying the parameters of GridConfig. The main parameters are: on_target (the behavior of an agent on the target, e.g., restart for LifeLong MAPF and nothing for classical MAPF), seed – to preserve the same generation of the map; agent; and their targets for scenario, size – used for cases without custom maps to specify the size of the map, density – the density of obstacles, num_agents – the number of agents, obs_radius – observation radius, collision_system – controls how conflicts are handled in the environment (we used a soft collision system for all of our experiments). The example of creation such instance is presented in Figure 11.
from pogema import pogema_v0, GridConfig, AnimationMonitor grid = """ .....#..... .....#..... ........... .....#..... .....#..... #.####..... .....###.## .....#..... .....#..... ........... .....#..... """ # Define new configuration with 6 randomly placed agents grid_config = GridConfig(map=grid, num_agents=6) # Create custom Pogema environment with AnimationMonitor env = AnimationMonitor(pogema_v0(grid_config=grid_config)) env.reset() # Saving SVG animation env.save_animation('four-rooms.svg') Refer to caption
Figure 11: Setting up a POGEMA instance with a custom map and generating an animation.
Visualization of the agents is a crucial tool for debugging algorithms, visually comparing them, and presenting the results. Many existing MARL environments lack such tools, or have limited visualization functionality, e.g., requiring running the simulator to provide replays, or offering visualizations only in one format (such as videos). In the POGEMA environment, there are three types of visualization formats. The first one is console rendering, which can be used with the default render methods of the environment; this approach is useful for local or server-side debugging. The preferred second option is SVG animations. An example of generating such a visualization is presented in the listing above. This approach allows displaying the results using any modern web browser. It provides the ability to highlight high-quality static images (e.g., as the images provided in the paper) or to display results on a website (e. g., animations of the POGEMA repository on GitHub). This format ensures high-quality vector graphics. The third option is to render the results to video format, which is useful for presentations and videos.
@@ Appendix E POGEMA Toolbox
The first one is registries to handle custom maps and algorithms. To create a custom map, the user first needs to define it using ASCII symbols or by uploading it from a file, and then register it using the toolbox (see Figure 13). The same approach is used to register and create algorithms (see Figure 12). In that listing, the registration of a simple algorithm is presented, which must include two methods: act and reset_states. This approach can also accommodate a set of hyperparameters which the Toolbox handles.
from pogema import BatchAStarAgent # Registring A* algorithm ToolboxRegistry.register_algorithm('A*', BatchAStarAgent) # Creating algorithm algo = ToolboxRegistry.create_algorithm("A*")
Figure 12: Example of registering the A* algorithm as an approach in the POGEMA Toolbox.
from pogema_toolbox.registry import ToolboxRegistry # Creating custom map custom_map = """ .......#. ...#...#. .#.###.#. """ # Registring custom map ToolboxRegistry.register_maps({"custom_map": custom_map})
Figure 13: Example of registering a custom map to the POGEMA Toolbox.
environment: # Configuring Test Environments name: Environment on_target: 'restart' max_episode_steps: 128 observation_type: 'POMAPF' collision_system: 'soft' seed: grid_search: [ 0, 1, 2, 3, 4, 5, 6, 7, 8, 9 ] num_agents: grid_search: [ 8, 16, 24, 32, 48, 64 ] map_name: grid_search: [ validation-mazes-000, validation-mazes-001, validation-mazes-002, validation-mazes-003, validation-mazes-004, validation-mazes-005, ] algorithms: # Specifying algorithms and it's hyperparameters RHCR_5_10: name: RHCR parallel_backend: 'balanced_dask' num_process: 32 simulation_window: 5 planning_window: 10 time_limit: 10 low_level_planner: 'SIPP' solver: 'PBS' results_views: # Defining results visualization MazesPlot: type: plot x: num_agents y: avg_throughput width: 4.0 height: 3.1 line_width: 2 use_log_scale_x: True legend_font_size: 8 font_size: 8 name: Mazes ticks: [8, 16, 24, 32, 48, 64] TabularThroughput: type: tabular print_results: True
Figure 14: Example of the POGEMA Toolbox configuration for parallel testing of the RHCR approach and visualization of its results.
Second, it provides a unified way of conducting distributed testing using Dask  [^16] https://github.com/dask/dask and defined configurations. An example of such a configuration is provided in Figure 14. The configuration is split into three main sections; the first one details the parameters of the POGEMA environment used for testing. It also includes iteration over the number of agents, seeds, and names of the map (which were registered beforehand). The unified grid_search tag allows for the examination of any existing parameter of the environment. The second part of the configurations is a list of algorithms to be tested. Each algorithm has its alias (which will be shown in the results) and name, which specifies the family of methods. It also includes a list of hyperparameters common to different approaches, e.g., the number of processes, parallel backend, etc., and the specific parameters of the algorithm.
@@ Appendix F Extended Related Work
Nevertheless, some of these drawbacks have already been addressed. SMAX Rutherford et al. (2023) provides a hardware-accelerated JAX version of the environment, but it cannot guarantee full compatibility since the StarCraft II engine is proprietary software. SMAC v2 Ellis et al. (2024) solves the problem of determinism, highlighting this issue in the original SMAC environments. Moreover, an evaluation protocol for the SMAC environment is proposed in Gorsane et al. (2022). Despite these efforts, it’s hard to say that these tasks require the generalization ability of the agent, since the training and evaluation are conducted on the same scenario.
@@
Multi-agent MuJoCo — In MAMuJoCo, the standard tasks involve agents controlling different sets of joints (or a single joint) within a simulated robot. This set of environments is a natural adaptation of the environment presented in the well-known MuJoCo physics engine Todorov et al. (2012). These tasks don’t require high generalization abilities or navigation. In the newer version, MuJoCo provides a hardware-accelerated version, forming the basis for Multi-agent BRAX Rutherford et al. (2023), which enhances performance and efficiency.
@@
Google Research Football — Google Research Football Kurach et al. (2020) is a multi-agent football simulator that provides a framework for cooperative or competitive multi-agent tasks. Despite the large number of possible scenarios in the football academy and the requirement for simple navigation, the tasks are highly specific to the studied domain. Additionally, the number of possible agents is limited. Moreover, the framework offers low scalability, requiring a heavy engine.
@@
Multi-robot warehouse — The multi-robot warehouse environment RWARE Papoudakis et al. (2021) simulates a warehouse with robots delivering requested goods. The environment is highly specific to delivery tasks; however, it doesn’t support procedurally generated scenarios, thus not requiring generalization abilities or an evaluation protocol. The best-performing solution Christianos et al. (2020) in this environment is trained on only 4 agents. The benchmark is highly related to multi-agent pathfinding tasks; however, it doesn’t provide centralized solution integration, which could serve both as an upper bound for learnable decentralized methods and as a source of expert demonstrations.
@@
Level-Based Foraging — Multi-agent environment LBF Papoudakis et al. (2021) simulates food collection by several autonomously navigating agents in a grid world. Each agent is assigned a level. Food is also randomly scattered, each having a level on its own. The collection of food is successful only if the sum of the levels of the agents involved in loading is equal to or higher than the level of the food. The agents are getting rewarded by level of food they collected normalized by their level and overall food level of the episode. The game requires cooperation but also the agents can emerge competitive behavior. The environment is very efficiently designed and very simple to set up; however, it doesn’t support procedurally generated scenarios, thus not requiring generalization abilities or an evaluation protocol.
@@
Flatland — The Flatland environment Mohanty et al. (2020) is designed to address the specific problem of fast, conflict-free train scheduling on a fixed railway map. This environment was created for the Flatland Competition Laurent et al. (2021). The overall task is centralized with full observation; however, there is an adaptation to partial observability for RL agents. Unfortunately, during several competitions, despite the presence of stochastic events, centralized solutions Li et al. (2021a) from operations research field have outperformed RL solutions by a large margin in both quality and speed. The environment is procedurally generated, which requires high generalization abilities, and the benchmark provides an evaluation protocol. A significant drawback is the extremely slow speed of the environment, which highly restricts large-scale learning.
@@
Overcooked — The Overcooked is a benchmark environment for fully cooperative human-AI task performance, based on the widely popular video game (Carroll et al., 2019). In the game, agents control chefs tasked to cook some dishes. Due to possible complexity of the cooking process, involving multistep decision-making, it requires emergence of cooperative behavior between the agents.
@@
Griddly — This is a grid-based game engine (Bamford, 2021), allowing to make various and diverse grid-world scenarios. The environment is very performance efficient, being able to make thousands step per second. Moreover, there is test coverage and continuous integration support, allowing open-source development. The engine provides support for different observation setups and maintains state history, making it useful for search based methods.
@@
Multi-player game simulators — Despite the popularity of multi-player games, it’s a challenging problem to develop simulators of the games that could be used for research purposes. One of the most popular adaptations are MineCraft MALMO (Johnson et al., 2016) that allows to utilize MineCraft as a configurable research platform for multi-agent research and model various agents’ interactions. In spite of the game’s flexible functionality, it depends on external runtime, so might be very hard to set up or extremely slow to iterate with. That’s why there are several alternatives that prioritize fast iteration over the environment complexity, like Neural MMO Suarez et al. (2024) that models a simple MMO RPG with agents with a shared resource pool. On top of that, there are even faster implementation, targeting coordination or cooperation, like Hide-and-Seek Baker et al. (2020), which models competition, or GoBigger Zhang et al. (2023), focusing on competition between cooperating populations.
Magent ––– Is a well-recognized environment within the community, though less widely used than SMAC. However, it lacks procedural generation capabilities, essential for testing agent generalization. The benchmark consists of six static scenarios, with the largest agent population observed in the Gather scenario, supporting up to 495 agents. The scenarios include Adversarial Pursuit, where red agents navigate obstacles to tag blue agents without causing damage; Battle, a large-scale team battle rewarding individual agent performance; Battlefield, similar to Battle but with fewer agents, but with predefined obstacles on the field; Combined Arms, a team battle featuring ranged and melee units with differing attack ranges, speeds, and health points; Gather, where agents compete for finite food resources requiring multiple “attacks” to consume; and Tiger-Deer, in which tigers team up to hunt deer, earning rewards through successful cooperation.
Gigastep — A GPU-accelerated multi-agent benchmark that supports both collaborative and adversarial tasks, continuous and discrete action spaces, and provides RGB image and feature vector observations. It features a diverse set of scenarios with heterogeneous agents, asymmetric teams, and varying objectives (e.g., tagging, waypoint following, hide-and-seek). However, it lacks procedural generation, which is crucial for testing generalization. Its evaluation protocol is minimal, relying on comparisons between PPO and hand-designed bots, offering limited insights into algorithmic diversity.
@@
Multi-agent Driving Simulators — Autonomous driving is one of the important practical applications of MARL, and Nocturne  Vinitsky et al. (2022) is a 2D simulator, written in C++, that focuses on different scenarios of interactions — e.g. intersections, roundabouts etc. The simulator is based on trajectories collected in real life, allowing it to model practical scenarios. This environment has evaluation protocols and supports open-source development with continuous integration and covered by tests. There are also environments, focusing on particular details of driving, for example, Multi-car Racing  Schwarting et al. (2021) that represents racing from bird’s eye view.
@@
Suits of multi-agent environments — These multi-agent environments are designed to be very simple benchmarks for specific tasks. Jumanji  Bonnet et al. (2023) is a set of environments for different multi-agent scenarios connected to combinatorial optimization and control, for example, routing or packing problems. With the purpose for each environment to be focused on the particular task, the overall suit doesn’t test generalization or enable procedural generation. Multi Particle Environments (MPE)  Lowe et al. (2017) is a communication oriented set of partially observable environments where particle agents are able to interact with fixed landmarks and each other, communicating with each other. SISL  Gupta et al. (2017) is a set of three dense reward environments was developed to have a simple benchmark for various cooperative scenarios. For environment suits, testing generalization, MeltingPot  Agapiou et al. (2022) comes into place. This set of the environments contains a diverse set of cooperative and general-sum partially observable games and maintains two populations of agents: focal (learning) and visiting (unknown to the environment) to benchmark generalization abilities of MARL algorithms. The set in based on the own game engine and might be extended quite easily.
@@
Real-world Engineering in Practice — Real-world engineering tasks can often be addressed by MARL solutions. IMP-MARL Leroy et al. (2024) provides a platform for evaluating the scalability of cooperative MARL methods responsible for planning inspections and repairs for specific system components, with the goal of minimizing maintenance costs. At the same time, agents must cooperate to minimize the overall risk of system failure. MATE Pan et al. (2022) addresses target coverage control challenges in real-world scenarios. It presents an asymmetric cooperative-competitive game featuring two groups of learning agents, cameras and targets, each with opposing goals.